From the window in a nearby building, the angle of elevation to the top of an apartment building is 28°. The angle of depression from the same window to the bottom of the apartment building is 34°. The window is 24 feet above the ground. How tall is the apartment building?
1
Expert's answer
2020-03-23T16:07:18-0400

As per the given question,

angle of elevation ="28^\circ"

angle of depression from the same window to the bottom of the building "=34^\circ"

The height of the window from the bottom(h) =24 feet

Let height of the building from the level of window is h'

For the depression from the window,

Let the distance between the building is x,

So, "\tan 34^\circ=\dfrac{h}{x}"

"x=\dfrac{24}{\tan 34^\circ}=35.58 feet"

Now, similarly,

For the above from the window,

"\tan 28^\circ=\dfrac{h'}{x}"

"\Rightarrow h'=x\tan 28^\circ feet"

"\Rightarrow h'=35.58\times tan(28^\circ)=18.92 feet"

Hence total height of the building = h+h'

"=24+18.92=42.92 feet"